# Verisimilitude as Boon and Bane: How People Initiate Opportunistic Interactions at Professional Events in Social VR

### v1

@@ 4.2. Capturing the Attention of Others
Clapping hands, making positive facial expressions, and reducing the proximity between avatars appeared as popular cues among many else to gain other’s attention in social VR. These popular cues share two qualities valued by our participants: they convey the social meaning of someone being “welcoming” or “receptive to others” with high verisimilitude, and they are visually appealing. Moreover, we found that participants did not make impulsive decisions about which specific cue(s) to activate for attention-grabbing at a given moment. Instead, they learned the constraints of each cue through trials-and-errors in social VR, using that insight to improve their initiation of opportunistic interactions with others. Below, we detail three specific constraints discussed by our participants, each affecting their use of certain cues over others (sections 4.2.1-4.2.3).
@@
We sought the other participants’ their reactions to the “virtual business cards” idea. While some felt this card exchange would really help “seal the deal and encourage more interaction to happen,” others disliked it. For them, one unique value of social VR is the flexibility built into interpersonal exchanges. People can make themselves “visible but also anonymous enough” in front of others. They can more easily avoid the social pressure of “feeling obligated to respond to others,” which is common in face-to-face interactions. However, the etiquette of card exchange, if introduced into the virtual environment, might impose too much social pressure on the recipient. It would not allow an equal level of flexibility to that existing in the exchange of other attention-grabbing cues, such as facial expressions via emojis and reduced proximity between avatars.
Together, our data shows that hand gestures, facial expressions through emojis, and proximity adjustment are all useful cues for people to capture others’ attentions in social VR. Participants discussed three constraints that could disrupt the verisimilitude of these cues. As our analysis has demonstrated, each cue was perceived as sensitive to some of these constraints but resilient to others. To mitigate the potential loss of social meanings, P6 reported that he followed a “give and take protocol:” when there was a person he really wished to speak with, she would first observe the attention-grabbing cue that person used most frequently, then he adopted the same cue to initiate his interaction with them. Notably, nobody else reported using the same strategy, probably due to the effort required to implement it. Issuing multiple cues simultaneously appeared as a much more common strategy used by most participants to gain other’s attention for opportunistic interaction.
@@ 4.4. Acquiring Established Interaction Protocols in social VR
Participants shared anecdotes witnessed from past events, where fully fluent social VR users could replicate their behaviors exactly as they would in physical reality. For example, those users could demonstrate natural body wiggling via their avatars. They sometimes even chose to leave virtual conference rooms by walking outside and then closing the virtual door, showcasing a sophisticated understanding of the hand controller mechanics as well as the virtual platform’s specific features. Participants commented that such “super users” might always be adept at overcoming constraints in VR and recovering the verisimilitude of all meaningful cues. However, none of our participants identified themselves as such. Even for participants with extensive experience hosting professional events in VR, they were confident in their “VR fluency” on the platforms they regularly used, but also highly aware that the technical protocol guiding people’s action possibilities could vary significantly across platforms.
@@ 5. Discussion
To recap, our work illustrates that, at professional events in social VR, initiating opportunistic interactions is challenging and should not be regarded as a singular process as in real-world settings. Rather, it comprises three successive steps. Each step features the intuitive adoption of a selected set of non-verbal cues, and it requires a progressively higher degree of verisimilitude than the preceding one(s) to distill actionable meanings from these cues (see Figure 1 for a summary). The technical infrastructure of today’s social VR platforms, unfortunately, often fails to support non-verbal exchanges with such subtleties. In this section, we elaborate on the empirical contributions as well as the design implications of the current research. Our reflections center around three primary takeaways derived from our findings: the challenges in estimating verisimilitude for initiating opportunistic interactions (Section 5.1), the unique role served by each stakeholder in initiating opportunist interactions (Section 5.2), and the preparation work that can be done in advance to ready people for initiating opportunistic interactions (Section 5.3).
@@ 5.1. Verisimilitude and Challenges in Estimating It
Figure 2. Three possible ways to indicate a VR user’s availability or willingness to engage in conversation: a) Status indicators as floating toggle icons above the avatar. b) Bubbles as visual cues that clearly indicate chat availability. c) Fading out avatars that are currently unavailable for social interaction. Illustrations generated by DALL-E 3 based on a user-provided prompt.
@@ 5.2. Social Responsibilities Delegated to the Attendee, the Host, or the System
Figure 3. Hosts and non-human embodied agents can help facilitate opportunistic conversations. a) VR platforms may notify hosts of potential hotspots where such conversations are likely to emerge. b) Embodied conversational agents (shown in red) can recommend icebreakers based on context, event topics, or user profiles. Illustrations generated by DALL-E 3 based on a user-provided prompt.
@@ 5.3. In Preparation for Successful Opportunistic Interactions at Social VR Events
We suspect that the lack of a unified policy in current social VR platform design is not due to feasibility concerns; rather, it is a deliberate choice shaped by dominant use cases in the past. As discussed earlier, VR has long been considered a medium for exploring creative forms of casual encounters (e.g., (Freeman and Acena, 2021)) and for gaming and social grooming among young generations (e.g., (Maloney et al., 2020a)). Much of the engaging and joyful user experience in these contexts is arguable tied to the unconventional and un-unified aspects of VR environment design. The recent adoption of social VR for professional events during and post the COVID pandemic pushes researchers and system builders to revisit design guideline for this medium, particularly in sticking the right balance between innovation and real-world resemblance for effective communication of non-verbal cues. Independent of the idea of making the VR environment more structured for users’ mental model formation, another approach is to facilitate people in distilling context-sensitive knowledge from the trial and error they will inevitably go through. Some of our research findings have shed light on this second approach. Specifically, participants in the current study recognized that system tutorials and user manual, both available via the platform they used, were valuable and legitimate resources to get themselves familiar with specific system features and the social functions being afforded. However, they preferred not to leverage these standalone resources, instead choosing to accumulate experiential knowledge through real interactions with others.

### v2

@@ 4.2. Capturing the Attention of Others
Clapping hands, making positive facial expressions, and reducing the proximity between avatars appeared as popular cues among many else to gain others’ attention in social VR. These popular cues share two qualities valued by our participants: they convey the social meaning of someone being “welcoming” or “receptive to others” with high verisimilitude, and they are visually appealing. Moreover, we found that participants did not make impulsive decisions about which specific cue(s) to activate for attention-grabbing at a given moment. Instead, they learned the constraints of each cue through trials-and-errors in social VR, using that insight to improve their initiation of opportunistic interactions with others. Below, we detail three specific constraints discussed by our participants, each affecting their use of certain cues over others (sections 4.2.1-4.2.3).
@@
We sought the other participants’ reactions to the “virtual business cards” idea. While some felt this card exchange would really help “seal the deal and encourage more interaction to happen,” others disliked it. For them, one unique value of social VR is the flexibility built into interpersonal exchanges. People can make themselves “visible but also anonymous enough” in front of others. They can more easily avoid the social pressure of “feeling obligated to respond to others,” which is common in face-to-face interactions. However, the etiquette of card exchange, if introduced into the virtual environment, might impose too much social pressure on the recipient. It would not allow an equal level of flexibility to that existing in the exchange of other attention-grabbing cues, such as facial expressions via emojis and reduced proximity between avatars.
Together, our data shows that hand gestures, facial expressions through emojis, and proximity adjustment are all useful cues for people to capture others’ attentions in social VR. Participants discussed three constraints that could disrupt the verisimilitude of these cues. As our analysis has demonstrated, each cue was perceived as sensitive to some of these constraints but resilient to others. To mitigate the potential loss of social meanings, P6 reported that he followed a “give and take protocol:” when there was a person he really wished to speak with, she would first observe the attention-grabbing cue that person used most frequently, then he adopted the same cue to initiate his interaction with them. Notably, nobody else reported using the same strategy, probably due to the effort required to implement it. Issuing multiple cues simultaneously appeared as a much more common strategy used by most participants to gain others’ attention for opportunistic interaction.
@@ 4.4. Acquiring Established Interaction Protocols in Social VR
Participants shared anecdotes witnessed from past events, where fully fluent social VR users could replicate their behaviors exactly as they would in physical reality. For example, those users could demonstrate natural body wiggling via their avatars. They sometimes even chose to leave virtual conference rooms by walking outside and then closing the virtual door, showcasing a sophisticated understanding of the hand controller mechanics as well as the virtual platform’s specific features. Participants commented that such “super-users” might always be adept at overcoming constraints in VR and recovering the verisimilitude of all meaningful cues. However, none of our participants identified themselves as such. Even for participants with extensive experience hosting professional events in VR, they were confident in their “VR fluency” on the platforms they regularly used, but also highly aware that the technical protocol guiding people’s action possibilities could vary significantly across platforms.
@@ 5. Discussion
To recap, our work illustrates that, at professional events in social VR, initiating opportunistic interactions is challenging and should not be regarded as a singular process as in real-world settings. Rather, it comprises three successive steps. Each step features the intuitive adoption of a selected set of non-verbal cues, and it requires a progressively higher degree of verisimilitude than the preceding one(s) to distill actionable meanings from these cues (see Figure 1 for a summary). The technical infrastructure of today’s social VR platforms, unfortunately, often fails to support non-verbal exchanges with such subtleties. In this section, we elaborate on the empirical contributions as well as the design implications of the current research. Our reflections center around three primary takeaways derived from our findings: the challenges in estimating verisimilitude for initiating opportunistic interactions (Section 5.1), the unique role served by each stakeholder in initiating opportunistic interactions (Section 5.2), and the preparation work that can be done in advance to ready people for initiating opportunistic interactions (Section 5.3).
@@ 5.1. Verisimilitude and Challenges in Estimating It
Figure 2. Three possible ways to indicate a VR user’s availability or willingness to engage in conversation: a) status indicators as floating toggle icons above the avatar; b) bubbles as visual cues that clearly indicate chat availability; c) fading out avatars that are currently unavailable for social interaction. Illustrations generated by DALL-E 3 based on a user-provided prompt.
@@ 5.2. Social Responsibilities Delegated to the Attendee, the Host, or the System
Figure 3. Hosts and non-human embodied agents can help facilitate opportunistic conversations. a) VR platforms may notify hosts (shown in green) of potential hotspots where such conversations are likely to emerge. b) Embodied conversational agents (shown in red) can recommend icebreakers based on context, event topics, or user profiles. Illustrations generated by DALL-E 3 based on a user-provided prompt.
@@ 5.3. In Preparation for Successful Opportunistic Interactions at Social VR Events
We suspect that the lack of a unified policy in current social VR platform design is not due to feasibility concerns; rather, it is a deliberate choice shaped by dominant use cases in the past. As discussed earlier, VR has long been considered a medium for exploring creative forms of casual encounters (e.g., (Freeman and Acena, 2021)) and for gaming and social grooming among young generations (e.g., (Maloney et al., 2020a)). Much of the engaging and joyful user experience in these contexts is arguable tied to the unconventional and un-unified aspects of VR environment design. The recent adoption of social VR for professional events during and post the COVID pandemic pushes researchers and system builders to revisit design guideline for this medium, particularly in striking the right balance between innovation and real-world resemblance for effective communication of non-verbal cues. Independent of the idea of making the VR environment more structured for users’ mental model formation, another approach is to facilitate people in distilling context-sensitive knowledge from the trial and error they will inevitably go through. Some of our research findings have shed light on this second approach. Specifically, participants in the current study recognized that system tutorials and user manuals, both available via the platform they used, were valuable and legitimate resources to get themselves familiar with specific system features and the social functions being afforded. However, they preferred not to leverage these standalone resources, instead choosing to accumulate experiential knowledge through real interactions with others.
@@ 8. Conclusion
Acknowledgements.
We are grateful to our participants for their time and insights, which made this work possible. We thank Yongle Zhang, Yimin Xiao, Cartor Hancock, and Yuhang Zhou, as well as the anonymous reviewers, for their valuable feedback on various versions of our paper. We also thank Hao-Chuan Wang, John Tang, and Jessica Vitak for earlier discussions that informed this work.